Grade upto college level Physical Chemistry

The electron energy in hydrogen atom is given by E = (-2.17 * 10=12)/n2 ergs. Calculate the energy required to remove an electron completely from the n = 2 orbit. What is the longest wavelength (in cm) of light that can be used to cause this transition?

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
To calculate the energy required to remove electron from atom, n = ∞ is to be taken.
Energy of an electron in nth orbit of hydrogen is given by
E = 21.7 * 10-12 * 1/n2 ergs
∴ ∆E = 21.7 * 10-12 (1/22 – 1/∞2)
= - 21.7 * 10-12 (1/4 – 0) = 21.7 * 10-12 * 1/4
= - 5.42 * 10-12 ergs
∴ ∆E = hc/λ (∵ v = c/λ)
Or λ = hc/∆E
Substituting the values, λ = 6.627 *10-27 *3 *1010/5.42 * 10-12
= 3.67 * 10-5 cm